Aggrastat at 15.2 mg in 72 mL is to be infused at 8 mcg/kg/hr in a patient who weighs 38 kg. At what flow rate in mL/hr will you set the pump?
Answer #1

The weight of the patient is 38 kgs

Per hour = 8 mcg \times 38 kgs

               = 304 mcg to be infused in an hour

convert mcg to mg = 0.3 mg per hour

Flow rate = order per hour \div available dose \times available volume

               = 0.3 \div 15.2 \times 72

               = 1.4 ml / hour